OpenStudy (anonymous):

When you simplify each side of the equation 5n – 2(3 – n) = 4(2n – 3), you will get the equivalent equation 7n – 6 = 8n - 12. True False

OpenStudy (anonymous):

false

OpenStudy (diyadiya):

5n – 2(3 – n) = 4(2n – 3) On the LHS -2(3-n)=-6+2n (distribute) 5n-6+2n=4(2n-3) 4(2n-3)=8n-12 5n-6+2n=8n-12 What do you think now?

OpenStudy (anonymous):

it is true because answer is 7n-6=8n-12